OpenStudy (anonymous):

Elliot has been running a lawn care business since 2000. He cuts grass, trims, and weed whacks yards for his customers throughout the season. Each year, he has increased his fee by the same amount. The table shows what Elliot charged each customer for two given years of his business: Year Lawn Care Fee 2000 $750 2010 $1350 A. What is the rate of change and initial value for Elliot’s business? How do you know? B. Write an equation in slope-intercept form to represent the fees that Elliot charges each year.

OpenStudy (anonymous):

ill give a metal to the best answer....

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

can you help real quick

OpenStudy (trojanpoem):

Initial value = value when he started ( in 2000) =$ 750 Each year he add a constants to that value let y be the amount he get after t years a amount increase each year y = 750 + at you're given that in 2010 his free = 1350 t = 2010 - 2000 = 10 year y = 1350 y = 750 + at 1350 = 750 + 10 a 600 = 10a a = 60 y = 750 + 60t ( his fee as a function of time) In slope form y = 60t + 750 so he increases his fee by 60 $ a year

OpenStudy (anonymous):

Thanks.....but how would i write it in paragraph form lol

OpenStudy (anonymous):

because i dont know how to get all those symbols in words and then word it correctly

OpenStudy (trojanpoem):

the rate of change : 60 dollar per year initial value for Elliot’s business: $750 How do you know ? -,- see my equations below. xD, maybe I am not good in writing paragraphs.

OpenStudy (trojanpoem):

maybe you could write : he started in 2000 and at that time his fee was 750 which is the initial value by 2010 his fee became 1350 so his fee increased each year by 50 $

OpenStudy (anonymous):

thank you muchly lol^_^

OpenStudy (trojanpoem):

You're welcome.